# ONEBench to Test Them All: Sample-Level Benchmarking Over Open-Ended Capabilities

### v1

@@ 3.1 Theoretical Foundations
We begin by postulating a ground-truth statistical model generating the data, which is converted into ordinal comparisons ($\mathcal{S}$) [^2] This contrasts with Zhang and Hardt [174], who view aggregation as classical voting, analysing tradeoffs in aggregating voter preferences rather than uncover an underlying ranking.. Specifically, we use a random-utility model [149], where each model $f_{i}$ is associated with a utility distribution $\mathcal{U}_{f_{i}}$. Preferences between models $f_{i}$ and $f_{j}$ are based on comparing sampled utilities, i.e., $f_{i}\prec f_{j}:=u(f_{i})<u(f_{j})$, where $u_{f}\sim\mathcal{U}_{f}$. Since computing maximum likelihood estimates over general random-utility models is computationally hard [163], we focus on the Plackett–Luce model [119, 94], the only known exception that allows for tractable maximum likelihood estimates (MLE).
@@
Property 1: Identifiability. We first ask: Are the utility distributions for all models recoverable? The Plackett-Luce model allows identifying the utility distribution (up to an arbitrary additive constant) if all models are compared via a directed path [57, 163] [^3] Recall that using the reference model $f_{\text{base}}$ removes the additive ambiguity.. Consistency and asymptotic normality hold under specific assumptions about the comparison graph [50].
@@
Property 2: Sample-Efficient Convergence from Sparse Data. Given that identifiability is asymptotic, we then ask: How sample-efficient is the algorithm for recovering the utility distribution? With partial rankings of size $k$, the MLE is surprisingly sample efficient while being minmax-optimal [46, 102]. Specifically, sampling $k$ model comparisons from the model set $|\mathcal{F}|$ independently and uniformly at random for $|\mathcal{D}|$ samples induces an expander graph with high probability, which provides guarantees on sample-efficiency of recovery, with $|\mathcal{D}|=\nicefrac{{\Omega(|\mathcal{F}|)}}{{k}}$ samples being necessary, and $|\mathcal{D}|=\nicefrac{{\Omega(|\mathcal{F}|\log|\mathcal{F}|)}}{{k}}$ samples being sufficient. Efficient algorithms like those in Agarwal et al. [1] and Maystre and Grossglauser [102] achieve these bounds. Rank-breaking techniques, used in our empirical evaluation, also offer near-optimal solutions [141].
Property 3: Active Aggregation. In ONEBench, we can strategically select model comparisons, by framing selection as an online multi-armed bandit problem. Sample efficiency can be improved with PAC guarantees [145, 124, 126], significantly outperforming random comparisons [103].
@@
Property 4: Social Properties. The Plackett-Luce model ensures computational efficiency and recoverability of the underlying ranking. However, to design democratic systems for decision-making, it is essential also to have fair aggregation. Ensuring fairness involves trade-offs[174], because different notions of fairness often conflict, and, depending on the intended application areas, differing or even opposing preferences may be valid [6, 10, 38]. The Plackett-Luce model offers “procedural fairness” [84] (Section 2.2), satisfying the following three criteria:
@@
  • Anonymity. All voters (samples) are treated equally, ensuring the system does not rely on a single vote. The rankings remain unchanged if the input sample set is permuted.
@@
  • Neutrality. The ranking is invariant to the identities of the models, ensuring fairness among alternatives. This means permuting the models similarly permutes the new ranking.
@@
  • Independence from Irrelevant Alternatives (IIA). The relative ranking of two models is unaffected by other alternatives in a given sample, as guaranteed by Luce’s axiom of choice [93]. This provides grounding for incomplete model evaluations.
@@ 3.2 Translating Theory to Practice
Figure 2: Top-10 model ranking changes across different aggregation methods. Our Plackett-Luce-based rank aggregation method (Ours) shows the most similarity to the Ground Truth model rankings (GT). However, there is a progressive degradation in ranking accuracy for LMArena (LMArena) and Elo (ELO). Comparisons are shown for ONEBench-LLM (top) and ONEBench-LMM (bottom). Our method best preserves the ranking of the top-10 models.
We now empirically validate our framework, aiming to show that: (i) the Plackett-Luce model works well on real-world data, (ii) our aggregation method is sample-efficient, and (iii) it handles high levels of incompleteness. Below, we describe our setup and address these points.
@@ 3.2.1 Setup
Benchmarks. We conduct experiments using four popular benchmarks with established model rankings based on average scores specific to the benchmark: HELM [79] and Open LLM Leaderboard [8] for LLMs, and VHELM [23] and LMMs-Eval [176] for LMMs. We use the score-based rankings over individual datasets as ground truth and define our sample pool as the sum of the data pools of all consituent datasets of a given benchmark. To test the faithfulness of our aggregation strategy we compare the resulting rankings to the original leaderboards. These leaderboards evaluate foundation models across varied tasks with different metrics, serving as good indicators of real-world performance.
@@
Ground Truth. The current system of benchmarking involves evaluating models on individual test sets and measuring the mean score per model. This holds even for benchmarks that combine test sets such as HELM [79], VHELM [23], etc. We consider these scores as the ground truth measurement and generate a ground truth model ranking from these scores. Since we aggregate benchmarks spanning multiple measurement metrics, we implement a min-max normalization of numeric measurements to bring all benchmark samples to the same 0-1 score range. Our final ground truth refers to the model rankings derived from the mean score across all benchmarks.
Methods. We evaluate three model ranking methods:
@@
(i) Elo Score [30]: A competitive game rating system adapted to rank models through pairwise comparisons, adjusting scores based on wins or losses to reflect win-rate reliability.
@@
(ii) LMArena Ranking [18]: A method for LLM ranking using the Bradley-Terry model [15], which estimates model rankings through Maximum Likelihood Estimation (MLE) based on pairwise comparisons using an underlying ELO model.
@@
(iii) Our Method: Our approach leverages the Plackett-Luce model [102] to aggregate pairwise comparisons using partial rank breaking [141], which achieves a significant speed-up for rank estimation methods.
@@
Metrics. We compare the rankings generated by each method to the ground-truth from the leaderboards using Kendall’s $\tau$, a standard correlation metric for rankings. Each method is tested three times, and we report the mean and variance. We additionally check that the top-$k$ models are reliably recovered.
3.2.2 Is Plackett-Luce a Good Fit for Real-World Data?
Metric HELM Open LLM VHELM LMMs-Eval
Leaderboard
Elo Score 0.347 ± 0.132 0.213 ± 0.065 0.639 ± 0.024 0.332 ± 0.109
LMArena Ranking 0.852 ± 0.001 0.969 ± 0.000 0.697 ± 0.000 0.427 ± 0.000
Our Method 0.877 ± 0.001 0.997 ± 0.000 0.799 ± 0.000 0.641 ± 0.000
Table 1: Kendall’s $\tau$ correlations to ground-truth ranking for different aggregation algorithms. Results show improvements over ELO and LMArena rankings, with notable correlation boosts on ONEBench-LMM leaderboards, including LMMs-Eval (41.65%) and VHELM (14.63%).
Figure 3: Sample-efficient convergence (top) and robustness to sparsity. Kendall $\tau$ between ground-truth ranking and different ranking methods as random individual data samples are dropped (top) and model measurements are randomly removed (bottom). Methods typically remain robust to missing data, with Plackett-Luce consistently achieving higher correlation, even with 95% measurements missing.
@@ 3.2.2 Is Plackett-Luce a Good Fit for Real-World Data?
Q1. Is it a good fit? We assess whether the Plackett-Luce model performs well on large-scale benchmark data by comparing our aggregation algorithm’s rankings to the leaderboard rankings. As shown in Table 1, our algorithm achieves a high positive Kendall’s $\tau$, indicating strong alignment with the ground truth rankings.
@@
Q3. Are the top-k models preserved? For practitioners, the critical concern is whether the top models are ranked correctly. Figure 2 shows that our algorithm effectively preserves the top-10 model rankings compared to ground truth, while outperforming state-of-the-art methods in maintaining accurate top-k rankings.
@@
Conclusion. The Plackett-Luce model fits real-world data well, outperforming other methods in both overall Kendall’s $\tau$ and top-10 model rankings, proving its effectiveness for large-scale benchmarks. The underlying reason is that we avoid the limitations of Elo-based methods, which rely on assumptions that do not apply to foundation models [13].
3.2.3 Sample Efficiency and Handling Incomplete Rankings
We now empirically test the sample efficiency and robustness to incomplete data of our framework.
@@ 3.2.3 Sample Efficiency and Handling Incomplete Rankings
Q1. Is Our Algorithm Sample-Efficient? We systematically reduce the number of samples and re-rank the models using various methods, calculating Kendall’s $\tau$ for each. Missing data is simulated from 0% to 99%, with 10% intervals until 90%, followed by 1% increments. As shown in Fig. 3, our method maintains stable performance even with up to 95% samples missing, demonstrating that it can achieve accurate rankings with far fewer data points—up to 20x less than current benchmarks.
@@
Q2. Can our Algorithm Aggregate Highly Sparse Rankings? We evaluate the method’s ability to handle highly incomplete data by removing model comparisons for each sample and re-ranking the models. We randomly remove a given fraction of model measurements from each sample and re-rank using the three aggregation methods. Again, we simulate data removal from 0% to 99%, with increments as before. As shown in Fig. 3, our method performs well even with 95% fewer model comparisons, proving it can recover accurate rankings with highly sparse data. This is crucial for ONEBench, where models cannot be expected to be evaluated on the entire data pool.
@@
Conclusion. Our algorithm provides significant sample efficiency, maintaining accurate rankings with 20x fewer data points, and is robust to highly sparse input rankings.
4 ONEBench: Creation & Capability Querying

### v2

@@ 3.1 Theoretical Foundations
We begin by postulating a ground-truth statistical model generating the data, which is converted into ordinal comparisons ($\mathcal{S}$). This is in contrast to Zhang and Hardt (2024), who view aggregation as classical voting and analyse tradeoffs in aggregating voter preferences rather than uncover an underlying ranking. Specifically, we use a random-utility model (Thurstone, 1927), where model $m_{i}$ is associated with utility distribution $\mathcal{U}_{m_{i}}$. Preferences between models $m_{i}$ and $m_{j}$ are based on comparing sampled utilities, i.e., $m_{i}\prec m_{j}:=u(m_{i})<u(m_{j})$, where $u_{m}\sim\mathcal{U}_{m}$. Since computing maximum likelihood estimates over general random-utility models is computationally hard (Xia, 2019), we focus on the Plackett–Luce model (Plackett, 1975; Luce, 1977), the only known exception that allows for tractable MLE.
@@
Property 1: Identifiability. We first ask: Are the utility distributions for all models recoverable? The Plackett-Luce model allows identifying the utility distribution (up to an arbitrary additive constant) if all models are compared via a directed path (Xia, 2019). Using reference model $m_{\text{base}}$ removes additive ambiguity. Consistency and asymptotic normality hold under specific assumptions about the comparison graph (Han and Xu, 2023).
@@
Property 2: Sample-Efficient Convergence from Sparse Data. Given that identifiability is asymptotic, we ask: How sample-efficient is the algorithm for recovering the utility distribution? With partial rankings of size $k$, the MLE is surprisingly sample efficient while being minmax-optimal (Maystre and Grossglauser, 2015). Sampling $k$ model comparisons from the model set $|\mathcal{M}|$ uniformly at random induces an expander graph with high probability, giving guarantees for sample-efficient recovery, with $\nicefrac{{\Omega(|\mathcal{M}|)}}{{k}}$ samples being necessary, and $\nicefrac{{\Omega(|\mathcal{M}|\log|\mathcal{M}|)}}{{k}}$ samples being sufficient. Efficient algorithms like Maystre and Grossglauser (2015) achieve these bounds. Rank-breaking techniques, used in our evaluation, offer near-optimal solutions (Soufiani et al., 2014).
@@
Property 3: Social Properties. The Plackett-Luce model offers computational efficiency and recoverability of the underlying ranking. However, designing democratic decision-making systems also requires fair aggregation. Ensuring fairness involves trade-offs (Zhang and Hardt, 2024), as different notions of fairness often conflict. Moreover, depending on the intended application areas, differing or even opposing preferences may be valid (Arrow, 1950). Plackett-Luce offers “procedural fairness" (List, 2022), satisfying:
@@
(i) Anonymity. All voters (samples) are treated equally, ensuring the system does not over-rely on any single vote. Rankings remain unchanged if the input sample set is permuted.
@@
(ii) Neutrality. The ranking is invariant to model identities, ensuring fairness among alternatives. This means permuting the models similarly permutes the resulting ranking.
@@
(iii) Independence from Irrelevant Alternatives. The relative ranking of two models is unaffected by other alternatives in a given sample, as guaranteed by Luce (1959). This provides grounding for incomplete model evaluations.
In Section 7, we discuss separability and pairwise majority consistency—two properties violated by the Plackett-Luce model.
@@ 3.2 Translating Theory to Practice
Figure 2: Top-10 model ranking changes across different aggregation methods. Plackett-Luce (Ours) shows the most similarity to the Ground Truth model rankings (GT). However, there is a progressive degradation in ranking accuracy for LMArena (LMArena) and Elo (ELO).
Here, we show that: (i) the Plackett-Luce model works well on real-world data, (ii) our aggregation method is sample-efficient, and (iii) it handles high levels of incompleteness. Below, we describe our setup and address these points.
@@ 3.2.1 Setup
Benchmarks. We conduct experiments using four popular benchmarks with established model rankings based on benchmark-specific average scores: HELM (Liang et al., 2023) and Open LLM Leaderboard (Beeching et al., 2023) for LLMs, and VHELM (CRFM, 2024) and LMMs-Eval (Zhang et al., 2024c) for LMMs. We define our data pool as the sum of all samples in the constituent datasets. To test the faithfulness of our aggregation strategy we compare the resulting rankings to the original leaderboards. These leaderboards evaluate models across varied tasks with different metrics, serving as good indicators of real-world performance.
@@
Ground Truth. The current system of benchmarking involves evaluating models on individual test sets and measuring the mean score per model. This holds even for benchmarks that combine test sets. We consider these scores as the ground truth measurement and generate a ground truth model ranking from these scores. Since we aggregate multiple measurement metrics, we implement a min-max normalization of numeric measurements to bring all benchmark samples to the same 0-1 score range. Our final ground truth refers to the model rankings derived from the mean score across all benchmarks. While this procedure is suboptimal, and indeed we propose to replace it with our method, there is no other ground truth available for real data. We show results on synthetic data in appendix D.
Methods. We evaluate three ranking methods:
@@
(i) Elo Score (Elo, 1967): A competitive game rating system adapted to rank models through pairwise comparisons, adjusting scores based on wins or losses to reflect win-rate reliability.
@@
(ii) LMArena Ranking: A ranking method based on the Bradley-Terry model (Bradley and Terry, 1952), using a Maximum Likelihood Estimation (MLE) based on pairwise comparisons with an underlying ELO model for rank aggregation.
@@
(iii) Ours: We leverage the Plackett-Luce model (Maystre and Grossglauser, 2015) to aggregate pairwise comparisons using partial rank breaking, speeding up rank estimation.
@@
Metrics. We compare the rankings generated by each method to the ground-truth from the leaderboards using Kendall’s $\tau$, a standard correlation metric for rankings. Each method is tested three times and we report the mean and variance. We also check that the top-$k$ models are reliably recovered.
Dataset Elo LMArena Ours
HELM 0.35 ± 0.13 0.85 ± 0.00 0.88 ± 0.00
Leaderboard 0.21 ± 0.07 0.97 ± 0.00 0.99 ± 0.00
VHELM 0.63 ± 0.02 0.69 ± 0.00 0.80 ± 0.00
LMMs-Eval 0.33 ± 0.11 0.42 ± 0.00 0.64 ± 0.00
Table 1: Comparison to other aggregation algorithms. Kendall’s $\boldsymbol{\tau}$ correlations to ground-truth ranking across individual benchmarks.
3.2.2 Plackett-Luce on Real-World Data
@@
Q1. Is it suitable? We evaluate the Plackett-Luce model on large-scale benchmarks by comparing the rankings produced by our aggregation algorithm to the leaderboard rankings. We achieve strong alignment with the ground truth rankings (Table 1).
@@
Q3. Are the top-k models preserved? A key concern for practitioners is whether the top models are ranked correctly. As shown in fig. 2, our algorithm reliably preserves the ground-truth top-10 rankings, while other aggregation methods show weak and inconsistent alignment.
@@
Conclusion. The Plackett-Luce model fits real-world data well, outperforming other methods in both overall Kendall rank correlation and top-10 rankings, proving its effectiveness for large-scale benchmarks. The underlying reason is that we avoid using Elo distributions, which rely on assumptions that do not apply to foundation models (Boubdir et al., 2023).
3.2.3 Sample Efficiency and Robustness
@@
Q1. Is Our Algorithm Sample-Efficient? We systematically reduce the number of data samples and re-rank the models using various methods, calculating Kendall’s $\tau$ for each. Missing data is simulated from 0% to 99%, with 10% intervals until 90%, followed by 1% increments. As shown in fig. 3, our method maintains stable performance even with up to 95% samples missing, demonstrating that it can achieve accurate rankings with up to 20 times fewer data points than current benchmarks.
Figure 3: Sample-efficient convergence and robustness to sparsity. Kendall rank correlation coefficient between ground-truth ranking and different ranking methods as we randomly remove individual data samples (top) and model measurements (bottom). Methods typically remain robust to missing data, with Plackett-Luce consistently achieving higher correlation, even with 95% measurements missing. The strongly convex Plackett-Luce model reliably converges to the same ranking.
@@
Q2. Can our Algorithm Aggregate Highly Sparse Rankings? We assess our method’s ability to handle incomplete data by randomly removing a fraction of model measurements from each sample and re-ranking using the three aggregation methods. We simulate data removal from 0% to 99%, with increments as before. As shown in fig. 3, our method remains effective even when 95% of model comparisons are missing, proving it can recover accurate rankings with highly sparse data. This is crucial for ONEBench, which is designed to continually expand to accommodate more models and datasets, where models cannot be expected to be evaluated on the entire data pool.
@@
Conclusion. Our method is sample efficient and robust to sparse input rankings, maintaining accurate rankings with 20 times fewer data points.
3.2.4 Comparison to other Social Choice Theory Frameworks
Following the evaluation protocol used in  Rofin et al. (2022), we compare our rank aggregation method with the Borda Count and the Dowdall system to provide a broader perspective on its performance against other social choice theory methods. As shown in table 2, our method generally outperforms all the social choice theory frameworks, with drastic boosts on the more heterogeneous and incomplete benchmarks (VHELM and LMMs-Eval). We attribute this to the fact that the Plackett-Luce method is a probabilistic scoring model that explicitly considers the uncertainty in rankings and captures inter-model agreement patterns, while Borda Count (by amplifying noise by assigning uniform linear weights to ranks) and the Dowdall System are both deterministic point-based methods.
Dataset Borda Dowdall Ours
HELM 0.81 ± 0.00 0.83 ± 0.00 0.88 ± 0.00
Leaderboard 0.95 ± 0.00 0.99 ± 0.00 0.99 ± 0.00
VHELM 0.35 ± 0.00 0.21 ± 0.00 0.79 ± 0.00
LMMs-Eval 0.08 ± 0.00 0.18 ± 0.00 0.64 ± 0.00
Table 2: Comparison to other social choice theory methods. Kendall rank correlations to ground-truth ranking across individual benchmarks.
4 ONEBench: Creation & Capability Probing
4.1 ONEBench-LLM
Data Pool $\mathcal{D}$. For ONEBench-LLM, we source data from the Open LLM Leaderboard, HELM, and LMArena. Open LLM Leaderboard and HELM aggregate several individual benchmarks, such as MMLU (Hendrycks et al., 2021a) and HellaSwag (Zellers et al., 2019), while LMArena uses pairwise model comparisons based on user-generated prompts. Metrics include F1 score, exact match (EM), and quasi-exact match (QEM), as well as pairwise preferences.
Models $\mathcal{M}$. For ONEBench-LLM, we use the 100 most downloaded models from Open LLM Leaderboard and all 79 models from HELM (as of v1.9.0), including both proprietary models like GPT-4o (OpenAI, 2024) and open-weights ones like LLaMA-3 (Meta, 2024).
4.2 ONEBench-LMM
Data Pool $\mathcal{D}$. For ONEBench-LMM, data is sourced from VHELM, LMMs-Eval, and WildVisionArena. Similar to ONEBench-LLM, VHELM and LMMs-Eval aggregate individual datasets like MMMU (Yue et al., 2024) and VQAv2 (Goyal et al., 2017), while WildVisionArena uses pairwise tests for LMMs through image-based chats. Measurements include binary metrics like EM, QEM, and real-valued scores like ROUGE (Lin, 2004). We augment pairwise comparisons from WildVisionArena with LLM-as-a-Judge preferences generated using Prometheus-2 (Kim et al., 2024), which correlate highly with human judgments.
Models $\mathcal{M}$. For ONEBench-LMM, we use 14 models from LMMs-Eval and 25 models from VHELM, including proprietary models like Gemini Pro Vision (Team et al., 2023) and open-weights models like LLaVA (Liu et al., 2023a).
4.3 Capability Probing
Given a query, the system retrieves relevant data samples using a combination of semantic and metadata search. This capability probing provides a personalized comparison of foundation models.
(i) Semantic search. We perform a nearest neighbours search in the embedding space of all-MiniLM-L6-v2 (Reimers and Gurevych, 2019) for language tasks and SigLIP-B16 (Zhai et al., 2023) for vision-language tasks, using cosine similarity. We retrieve top $k$ samples for a given concept with tuned thresholds of $0.3$ for ONEBench-LLM and $0.7$ for ONEBench-LMM.
(ii) Metadata search. We verify that per-sample metadata satisfies the constraints defined in the query. Some benchmarks, such as MMMU, are equipped with detailed metadata, including categories like image type, question type, area, etc.
Using these search mechanisms, we retrieve relevant samples from the data pool and aggregate ordinal model rankings with the Plackett-Luce model. We test ONEBench with a curated set of 50 concepts ranging from domain-specific knowledge, such as the Coriolis effect, to broader academic disciplines like neuroscience, and objects like the iPad (fig. 4 and appendix F).
Metric LLM LMM
Number of concepts 40 50
Cohen-$\kappa$ 0.79 0.91
mAP 0.85 0.73
CMC@1 0.95 0.94
CMC@10 1.00 0.96
Table 3: Capability Probing (Quantitative): summary of accuracy and retrieval metrics.
Insight 1. Retrieved samples are accurate and well-aligned with target concepts. Manual filtering out of incorrect samples by expert annotators [^2] Inter-annotator agreement in table 3 shows strong consistency between annotators. Annotators volunteered and provided informed consent. IRB approval was not obtained. confirms high retrieval precision, with a mean Average Precision (mAP) of 0.85 (LLM) and 0.73 (LMM). These results, shown in table 3, indicate that our system reliably retrieves samples matching the intended capabilities, with room for improvement on some underrepresented concepts. Please refer to the per-concept AP in table 10 for a better indicator. Note that the retrieval mechanism is expected to only improve with better models and larger test sets covering more diverse capabilities.
Insight 2. Top models vary significantly across queries. A key aspect of evaluating capability-specific performance is checking whether the set of top-performing models remains consistent across different queries. If rankings were largely stable, with the same models dominating regardless of the capability assessed, fine-grained querying would offer limited practical value. However, our results in fig. 4 and fig. 6 reveal substantial variation in the top-$k$ models across domains and concepts. This variation highlights that relying solely on global leaderboards can obscure meaningful differences. These findings confirm that ONEBench enables more targeted model selection by surfacing the most appropriate candidates for a given capability query.
Samples Retrieved Random $\tau$ Query $\tau$
100 0.78 ± 0.04 0.68 ± 0.10
1,000 0.93 ± 0.01 0.83 ± 0.05
10,000 0.98 ± 0.00 0.91 ± 0.03
Table 4: Capability probing vs. random retrieval: Average Kendall’s $\tau$ with the global ranking.
Insight 3. More specific queries lead to more distinct model rankings. We conduct an experiment analyzing the divergence between global and capability-specific model orderings to assess how domain specificity affects ranking. Similar rankings would suggest that domain-specific evaluation adds limited value. For each of the 40 concepts in appendix F, we retrieve $n$ relevant samples and compute the average Kendall rank correlation with the global ranking. As a baseline, we perform five random retrievals of $n$ samples from the full pool and compute their average correlation. Results in table 4 show that concept-specific rankings are consistently less correlated with the global ranking than random ones, even as $n$ increases. This suggests that models specialize across domains and should be evaluated accordingly.
5 Related Works
Recent multi-task benchmarks, such as GLUE (Wang et al., 2019b), SuperGLUE (Wang et al., 2019a), and BigBench (Srivastava et al., 2023), test the broad capabilities of foundation models. However, these benchmarks use arithmetic mean for task aggregation (Beeching et al., 2023) which can distort rankings (Zhang and Hardt, 2024) and is sensitive to outliers (Agarwal et al., 2021) or missing scores (Himmi et al., 2023). ONEBench addresses these by enabling sample reuse, avoiding task selection bias (Dominguez-Olmedo et al., 2024). Inspired by social choice theory, ONEBench employs ordinal rankings and the Plackett-Luce model for aggregation, which is robust to irrelevant alternatives and outliers. Moreover, ONEBench reduces evaluation costs, similar to compressed subsets (Zhao et al., 2024) and lifelong benchmarks (Prabhu et al., 2024). Further, by flexibly integrating diverse sample and measurements contributions, ONEBench can be more inclusive than traditional benchmarks dominated by well-funded institutions (Pouget et al., 2024; Nguyen et al., 2024). We provide an expanded review in appendix G.
6 Conclusions and Open Problems
We introduce ONEBench, an open-ended benchmarking framework for foundation models. Our open, democratized benchmarking methodology allows various stakeholders to contribute evaluation samples and model measurements with detailed metadata. This affords creating customized benchmarks and testing arbitrary capabilities with semantic and structured searches. We provide an aggregation mechanism that is both theoretically grounded and empirically validated to be robust to incomplete data and heterogeneous measurements across evaluations. We demonstrate the utility of ONEBench for LLMs and LMMs, showing how dynamic probing reveals new insights into model performance on specific tasks and concepts. This combination of theoretical rigour, empirical results, and practical flexibility makes ONEBench a valuable tool for comprehensive evaluation.
7 Limitations
Our approach, while promising, comes with its share of challenges. We highlight three key issues.
Effects of Aggregation. Combining different types of evaluation data into a single ranking risks oversimplifying important performance differences. We mitigate this by introducing flexible querying. Furthermore, conversion to pairwise ranking leads to loss of information which could hurt aggregation algorithms. However, in real-world scenarios pairwise measurements perform better, despite information loss (Shah et al., 2014).
Reproducibility. The dynamic nature of capability querying and the expanding sample pool, though useful, makes it harder to maintain consistency and can introduce bias during data collection and aggregation. However, our framework supports consistent evaluation, as each sample has a unique identifier and the querying mechanism is deterministic. Reproducibility can be achieved simply by specifying which samples are included.
Statistical Modeling Assumptions. Our reliance on statistical models like Plackett–Luce might make assumptions about data distribution that may not always hold, affecting the reliability of our results. This limitation is not unique to our approach, but holds for any method that makes modeling assumptions. Noothigattu et al. (2020) show that Plackett–Luce based models do not satisfy the following axioms:
(i) Pairwise Majority Consistency. If pairwise preference order across models are consistent: a > b, b > c and a > c, then ranking should preserve the consistency: a > b > c.
(ii) Separability. If model a is higher than model b in MLE estimate scores in two input sets, a must be higher than b in MLE estimate scores of their combined set. We show empirically that violating separability is not a problem in practice. We present results in appendix E.
Overall, we believe democratic, open-ended benchmarking is an impactful direction to explore, despite the apparent limitations.
8 Broad Impacts
Our work could have a meaningful impact on efficacy of benchmarking for foundation models. With ONEBench, we offer a benchmarking framework that can adapt to different domains, allowing for more inclusive and transparent evaluation practices, empowering researchers and downstream practitioners. By making benchmarking more accessible, we hope to encourage fairness, reproducibility, and innovation in how evaluation frameworks are designed. In the long run, this approach can help build a deeper understanding of foundation models across both language and vision–language tasks. We do not believe that there are any immediate negative societal consequences as a result of this work, but caution that all findings are preliminary and need additional evaluation before deployment.
Acknowledgements
The authors would like to thank (in alphabetic order): Nikhil Chandak, Karel D’Oosterlinck, Shashwat Goel, Shyamgopal Karthik, Bálint Mucsányi and Tim Weiland for their helpful suggestions and feedback. Special thanks to Bo Li for (promptly and consistently) providing sample-level model evaluation logs of LMMs-Eval, Yujie Lu for providing early access to the battle data of Vision Arena and Palzer Lama for creating the logo of ONEBench and helping with figures. The authors would also like to thank Federico D’Agostino for recommending the title of the paper and the name of the benchmark. AP and MB acknowledge financial support by the Federal Ministry of Education and Research (BMBF), FKZ: 011524085B and Open Philanthropy Foundation funded by the Good Ventures Foundation. VU and SD thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VU is supported by a Google PhD Fellowship in Machine Intelligence. MB acknowledges support via the CRC 1233 on Robust Vision and from the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
Appendix A Datasets used in ONEBench: Further Details